# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ 7 Conclusions
Figure 19: Combined constraint from crust thermal relaxation (Sec. 6.1), neutron star cooling (Sec. 6.2), and pulsar glitches (Sec. 6.3) in our three scenarios accounting for modification of nuclear interactions in $\theta=\pi$ matter. Neutron star cooling is the dominant signature for $\varepsilon\ll 1$ while crust thermal relaxation and pulsar glitches set the strongest constraints for $\varepsilon\sim 1$. We consider the pink shaded region (Scenario A, no modification to nuclear interactions) to be the conservative benchmark constraint. Increasingly attractive nuclear interactions in axion condensed matter (Scenarios B and C) would cause the modification of crust structure to occur closer to the QCD axion line. Parameter space as in Fig. 14.
@@
In this paper, we present a comprehensive study of the neutron star equation of state in the presence of exceptionally light QCD axions. In dense objects, the axion potential receives corrections proportional to the nucleon density due to axion couplings to QCD. For a wide range of $\varepsilon$ (see Eq. (38)), a neutron star is both dense enough such that the minimum of the axion potential shifts to $\theta=\pi$, and large enough such that it is energetically favorable for a $\theta=\pi$ region to emerge at the center of the star and an axion domain wall profile between the $\theta=\pi$ inside and $\theta=0$ outside to develop. In the $\theta=\pi$ region, there are three main effects, namely, a negative axionic pressure of order $-m_{a}^{2}f_{a}^{2}$, lighter nucleon masses, and a decrease of the pion mass from $138\,{\rm MeV}$ to about $80\,{\rm MeV}$. The latter can lead to modification of nuclear interactions.
@@
We study these changes with Chiral Effective Field Theory and Relativistic Mean Field Theory. The results of ChiEFT are somewhat inconclusive at $\theta=\pi$ and a pion mass of about $80\,{\rm MeV}$, with the convergence of many-body perturbation theory becoming worse as the pion mass is decreased, but an initial assessment indicates that the nuclear force may become significantly more attractive at $\theta=\pi$ leading to self-bound neutron matter and axion condensation at $n_{B}\simeq n_{\rm sat}$ even for $\varepsilon\simeq 1$. We model this reduction of the pion mass in RMFT with a decrease of the $\sigma$ mass ($K_{\sigma}>0$, see definition given by Eq. (21) and related discussion), resulting in an increase of the strength of the attractive part of the nuclear interaction. Using RMFT, we calculate modifications to the equation of state of nuclear matter at different densities at $\theta=\pi$. Combining these modifications to the nucleon masses and nuclear interaction strength, the negative axion pressure, axion domain wall profile, and gravity, we compute the neutron star equation of state and structure at different $m_{a}$ and $\varepsilon$ shown in Figs. 5, 7, and 8. Qualitatively, the negative axion pressure, the lighter nucleon masses and more attractive nuclear force at $\theta=\pi$ lead to neutron stars with a reduced or missing inner crust and a much reduced outer crust that is partially or entirely at a negative pressure in the axion domain wall for $\varepsilon\lesssim 0.1$. Our preliminary investigations of the $m_{\pi}-$dependence of the nuclear EOS in ChiEFT suggests the intriguing possibility of axion condensation for $\varepsilon\simeq 1$ at $n_{B}\lesssim 2n_{\rm sat}$ and will be explored separately in future work.
@@
This finding has observational consequences that are significantly constrained by isolated neutron star cooling (also recently considered in Ref. nscooling_qcdaxion2024 ), crust thermal relaxation, and pulsar glitches. The combination of these constraints is shown in Fig. 19. These all depend primarily on the equation of state at $n_{B}\lesssim n_{\rm sat}$ where our treatment of the interaction between axions and nucleons is robust. As the most conservative bound, Scenario A in the figure shall be considered as the benchmark constraint, while Scenarios B and C shall be considered as a generous estimate of the uncertainty and possibly stronger constraints to be drawn once the $\theta$-dependence of QCD has been explored in greater detail. In particular, the pion mass dependence of the low energy constants in ChiEFT must be determined from Lattice QCD and the convergence of nuclear many-body perturbation theory at decreased pion mass must be addressed. Additionally, in the domain wall region, parity-violating nuclear interactions may become relevant and should be included in a detailed treatment in ChiEFT. More details are available in App. A.
@@
We also suggest two intriguing observational consequences that require further observational and theoretical investigation. As shown in Fig. 9, the mass-radius relation of neutron stars with a $\theta=\pi$ core is indicative of rapid stiffening of the equation of state at intermediate density. Additionally, for some $\varepsilon$ there are missing crust layers, resulting in very small radii for low mass neutron stars, similar to strange quark stars. With data from the NICER x-ray observations as well as neutron star merger measurements with LVK and future gravitational wave observatories, a precisely known mass-radius relation would strongly constrain the presence or absence of a $\theta=\pi$ core. A second intriguing possibility is self-bound objects at $\theta=\pi$ which we call $\pi$-balls, with densities close to nuclear density and stabilized by the attractive axion potential rather than gravity. These objects can have different masses and sizes, ranging from lighter than typical asteroid masses to planet masses. Further studies are required to understand their formation in violent events such as supernovae or neutron star mergers, their growth in the interstellar and intergalactic medium, as well as the resulting observational signatures.
@@ Acknowledgments
The U.S. DOE supported the work of M. K. and S. R. under Grant No. DE-FG02- 00ER41132. S. R. also thanks the members of the N3AS Physics Frontier Center, funded by the NSF Grant No. PHY-2020275 for useful conversations. The work of M. B. was supported by the U. S. Department of Energy grant. M.B. is supported by the U.S. Department of Energy Office of Science under Award Number DE-SC0024375 and the Department of Physics and College of Arts and Science at the University of Washington. M. B. is also grateful for the hospitality of Perimeter Institute, where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work was also supported by a grant from the Simons Foundation (1034867, Dittrich). C. D. acknowledges support from the National Science Foundation under award PHY 2339043. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance award DE-SC0013617.
@@ Appendix A Additional nuclear interactions in the axion domain wall
The energy density and pressure of the axion domain wall has a somewhat weak effect on the EoS of typical neutron stars. However, inside the axion domain wall, the properties of nuclear matter can be affected by the following additional terms Ubaldi:2008nf ,
@@
where $\mathcal{N}$ contains both nucleon fields and $\tau_{3}$ is the third Pauli matrix in isospin space. This parity-violating mass term and pion-nucleon interaction only turn on in regions of finite $\theta$ between 00 and $\pi$. If we expect $c_{+}$ and $c_{4}$ to be of order unity, the interaction between nucleons and pions is smaller than the leading nucleon pion interaction strength by $m_{q}/f_{\pi}\simeq 10^{-1}$ and the parity violating mass is a correction of order $m_{q}/m_{N}\simeq 10^{-3}$. We do not expect these interactions to matter for sufficient large axion masses. However, they might lead to interesting signatures for small axion masses (thick domain wall), especially in the $\mathbb{Z}_{\mathcal{N}}$ model.
@@ Appendix B RMFT corrections to the domain wall calculation
where $n_{B}^{(s)}$ is the total scalar density of baryons. In the outer crust and core, the expression in terms of the mean field value of the $\sigma$ field is simple.

### v2

@@ VII Conclusions
Figure 19: Combined constraint from crust thermal relaxation (Sec. VI.1), neutron star cooling (Sec. VI.2), and pulsar glitches (Sec. VI.3) in our three scenarios accounting for modification of nuclear interactions in $\theta=\pi$ matter. Neutron star cooling is the dominant signature for $\varepsilon\ll 1$ while crust thermal relaxation and pulsar glitches set the strongest constraints for $\varepsilon\sim 1$. We consider the pink shaded region (Scenario A, no modification to nuclear interactions) to be the conservative benchmark constraint. Increasingly attractive nuclear interactions in axion condensed matter (Scenarios B and C) would cause the modification of crust structure to occur closer to the QCD axion line. Parameter space as in Fig. 14.
@@
In this paper, we present a comprehensive study of the neutron star equation of state in the presence of exceptionally light QCD axions. In dense objects, the axion potential receives corrections proportional to the nucleon density due to axion couplings to QCD. For a wide range of $\varepsilon$ (see Eq. (4)), a neutron star is both dense enough such that the minimum of the axion potential shifts to $\theta=\pi$, and large enough such that it is energetically favorable for a $\theta=\pi$ region to emerge at the center of the star and an axion domain wall profile between the $\theta=\pi$ inside and $\theta=0$ outside to develop. In the $\theta=\pi$ region, there are three main effects, namely, a negative axionic pressure of order $-m_{a}^{2}f_{a}^{2}$, lighter nucleon masses, and a decrease of the pion mass from $138\,{\rm MeV}$ to about $80\,{\rm MeV}$. The latter can lead to the modification of nuclear interactions.
@@
We study these changes with Chiral Effective Field Theory and Relativistic Mean Field Theory. The results of ChiEFT are somewhat inconclusive at $\theta=\pi$ and a pion mass of about $80\,{\rm MeV}$, with the convergence of many-body perturbation theory becoming worse as the pion mass is decreased. Nevertheless, an initial assessment indicates that the nuclear force may become significantly more attractive at $\theta=\pi$ leading to self-bound neutron matter and axion condensation at $n_{B}\simeq n_{\rm sat}$ even for $\varepsilon\simeq 1$. We model this reduction of the pion mass in RMFT with a decrease of the $\sigma$ mass ($K_{\sigma}>0$, see definition given by Eq. (21), table 1, and related discussion), resulting in an increase of the strength of the attractive part of the nuclear interaction. Using RMFT, we calculate modifications to the equation of state of nuclear matter at different densities at $\theta=\pi$. Combining these modifications to the nucleon masses and nuclear interaction strength, the negative axion pressure, axion domain wall profile, and gravity, we compute the neutron star equation of state and structure at different $m_{a}$ and $\varepsilon$ shown in Figs. 5, 7, and 8. Qualitatively, the negative axion pressure, the lighter nucleon masses, and the more attractive nuclear force at $\theta=\pi$ lead to neutron stars with a reduced or missing inner crust and a much reduced outer crust that is partially or entirely at a negative pressure in the axion domain wall for $\varepsilon\lesssim 0.1$. Our preliminary investigations of the $m_{\pi}-$dependence of the nuclear EOS in ChiEFT suggest the intriguing possibility of axion condensation for $\varepsilon\simeq 1$ at $n_{B}\lesssim 2n_{\rm sat}$ and will be explored separately in future work.
@@
This finding has observational consequences that are significantly constrained by isolated neutron star cooling (also recently considered in Ref. [40]), crust thermal relaxation, and pulsar glitches. The combination of these constraints is shown in Fig. 19. These all depend primarily on the equation of state at $n_{B}\lesssim n_{\rm sat}$ where our treatment of the interaction between axions and nucleons is robust. As the most conservative bound, Scenario A in the figure shall be considered as the benchmark constraint, while Scenarios B and C shall be considered as a generous estimate of the uncertainty and possibly stronger constraints to be drawn once the $\theta$-dependence of QCD has been explored in greater detail. In particular, the pion mass dependence of the low energy constants in ChiEFT must be determined from Lattice QCD and the convergence of nuclear many-body perturbation theory at decreased pion mass must be addressed. Additionally, in the domain wall region, parity-violating nuclear interactions may become relevant and should be included in a detailed treatment in ChiEFT. More details are available in App. A.
@@
We also suggest two intriguing observational consequences that require further observational and theoretical investigation, also previously noted in [45]. As shown in Fig. 9, the mass-radius relation of neutron stars with a $\theta=\pi$ core is indicative of rapid stiffening of the equation of state at intermediate density. Additionally, for some $\varepsilon$ there are missing crust layers, resulting in very small radii for low-mass neutron stars, similar to strange quark stars. With data from the NICER x-ray observations as well as neutron star merger measurements with LVK and future gravitational wave observatories, a precisely known mass-radius relation would strongly constrain the presence or absence of a $\theta=\pi$ core. A second intriguing possibility is self-bound objects at $\theta=\pi$ which we call $\pi$-balls, with densities close to nuclear density and stabilized by the attractive axion potential rather than gravity. These objects can have different masses and sizes, ranging from planet to white dwarf masses. Further studies are required to understand their formation in violent events such as supernovae or neutron star mergers, their growth in the interstellar and intergalactic medium, as well as the resulting observational signatures.
@@ Acknowledgments
We thank Stefan Stelzl for helpful discussions and correspondence, as well as Anson Hook and Ken McElvain for helpful discussions. The U.S. Department of Energy supported the work of M. K. and S. R. under Grant No. DE-FG02-00ER41132. S. R. also thanks the members of the N3AS Physics Frontier Center, funded by the NSF Grant No. PHY-2020275 for useful conversations. The work of M. B. was supported by the U.S. Department of Energy grant. M.B. is supported by the U.S. Department of Energy Office of Science under Award Number DE-SC0024375 and the Department of Physics and College of Arts and Science at the University of Washington. M. B. is also grateful for the hospitality of Perimeter Institute, where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work was also supported by a grant from the Simons Foundation (1034867, Dittrich). C. D. acknowledges support from the National Science Foundation under award PHY 2339043. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance award DE-SC0013617.
@@ Appendix A Additional nuclear interactions in the axion domain wall
The energy density and pressure of the axion domain wall has a somewhat weak effect on the EoS of typical neutron stars. However, inside the axion domain wall, the properties of nuclear matter can be affected by the following additional terms [55],
@@
where $\mathcal{N}$ contains both nucleon fields and $\tau_{3}$ is the third Pauli matrix in isospin space. This parity-violating mass term and pion-nucleon interaction only turn on in regions of finite $\theta$ between 00 and $\pi$. If we expect $c_{-}$, $c_{+}$, and $c_{4}$ to be of order unity, the interaction between nucleons and pions is smaller than the leading nucleon pion interaction strength by $m_{q}/f_{\pi}\simeq 10^{-1}$ and the parity-violating mass is a correction of order $m_{q}/m_{N}\simeq 10^{-3}$. We do not expect these interactions to matter for sufficiently large axion masses. However, they might lead to interesting signatures for small axion masses (thick domain wall), especially in the $\mathbb{Z}_{\mathcal{N}}$ model.
@@ Appendix B RMFT corrections to the domain wall calculation
where $n_{B}^{(s)}$ is the total scalar density of baryons. In the outer crust and core, the expression in terms of the mean-field value of the $\sigma$ field is simple.